# Gubkin str. 8, 119991 Moscow, Russia

### v1

@@ 5.2 Turning on cross-gradients: $\mathbf{Z}_{t}=\mathbf{1}$ and $\mathbf{Y}\neq 0$
so the kernel $\mathcal{K}$ is literally the (direction–independent) two–particle density of states per RG decade, while the weights (5) are obtained by dressing this phase space by the constant coherence factors in the projectors $\Pi_{\pm}$.
@@ 6 Conclusion
Technically, we have shown that in dimensional regularization with minimal subtraction near $D=4-\epsilon$ all one–loop ultraviolet divergences are governed by two universal master objects: the tadpole with a single mass insertion and the zero–momentum bubble. Using a spectral decomposition of the anisotropic quadratic operator, we derived compact, basis–covariant expressions for their pole parts in terms of the direction–dependent eigenvalues and projectors of the kinetic matrix. Both master residues are dressed by a single positive kernel, which admits a transparent interpretation as a two–particle energy denominator, or equivalently as a two–particle phase–space weight per RG decade. This representation makes the physical origin of all anisotropy–dependent coefficients in the beta functions explicit.
@@ A.1 Tadpole integral
with the use of (33). After introduction of such regulator into (43) the original integral
@@
where $\mathbf{p}\equiv\rho\hat{\mathbf{n}}$ with $\rho\equiv|\mathbf{p}|\geq 0$, and $x\equiv\rho^{2};\Omega_{d-1}$ is the solid angle on the unit sphere and $\Omega_{d-1}=2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$. We also remind, that $\langle\ldots\rangle_{\hat{\mathbf{n}}}$ is given by (46). Hence, we arrive to the regulated representation
@@
Our aim is to regularize the divergencies in the latter expression. To this end let us use Schwinger parametrization
@@
$\times\int_{0}^{\infty}t^{-1/2}e^{-t\delta_{\rm IR}^{2}}dt\;t^{-d/2}\,\sum_{a}\left\langle s_{a}(\hat{\mathbf{n}})^{-d/2}\,\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}(\hat{\mathbf{n}})\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}$. (165)
@@
$\mathbf{T}_{0}(\delta_{\rm IR})$ $=\frac{1}{2^{d+1}\pi^{d/2}\sqrt{\pi}}\int_{0}^{\infty}t^{-\frac{(d+1)}{2}}e^{-t\delta_{\rm IR}^{2}}dt\;\sum_{a}\left\langle s_{a}^{-\frac{d}{2}}\,\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}(\hat{\mathbf{n}})\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}$. (166)
where we use $\Omega_{d-1}=2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$. Using $D=4-\epsilon$ (so $d=3-\epsilon$), the $L=0$ contribution $\mathbf{T}_{0}$ is a scaleless power-divergent integral and vanishes in pure dimensional regularization. With the auxiliary IR mass $\delta_{\rm IR}>0$ kept finite one finds $\mathbf{T}_{0}(\delta_{\rm IR})\propto\delta_{\rm IR}^{2}$ (and may contain a pole $\propto\delta_{\rm IR}^{2}/\epsilon$), hence it vanishes as $\delta_{\rm IR}\to 0$ and does not contribute to the MS counterterms of the original theory [^7] I.e. in $d=3-\varepsilon$ we use $\int_{0}^{\infty}dtt^{-\frac{d+1}{2}}e^{-t\delta_{\mathrm{IR}}^{2}}=\left(\delta_{\mathrm{IR}}^{2}\right)^{\frac{d-1}{2}}\Gamma\left(\frac{1-d}{2}\right)=\left(\delta_{\mathrm{IR}}^{2}\right)^{1-\varepsilon/2}\Gamma\left(-1+\frac{\varepsilon}{2}\right)\sim-\frac{2}{\varepsilon}\delta_{\mathrm{IR}}^{2}+\mathcal{O}\left(\delta_{\mathrm{IR}}^{2}\ln\delta_{\mathrm{IR}}^{2}\right)$. Therefore the MS pole relevant for renormalization starts at $\mathcal{O}(\mathbf{L})$, i.e. comes from $\mathbf{T}_{1}$.
@@ A.2 Bubble integrals
$\mathbb{B}_{ij;kl}(k)=\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\mathbf{G}_{ik}(p)\mathbf{G}_{jl}(p+k)$
@@
$\mathbb{B}_{ij;kl}(0;\delta_{\rm IR})\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\,\mathbf{G}_{\delta,ik}(p)\,\mathbf{G}_{\delta,jl}(p),$ (177)
@@
$\mathbb{B}_{ij;kl}(0;\delta_{\rm IR})=\frac{1}{2}\int\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\sum_{a,b}\tilde{\mathbf{P}}_{a}(\mathbf{p})_{ik}\,\tilde{\mathbf{P}}_{b}(\mathbf{p})_{jl}\,\mathcal{K}\!\big(\kappa_{a}(\mathbf{p})+\delta_{\rm IR}^{2},\;\kappa_{b}(\mathbf{p})+\delta_{\rm IR}^{2}\big),$ (181)
@@
Switchning to $x\equiv|\mathbf{p}|^{2}$ and use $d^{d}\mathbf{p}=\frac{\Omega_{d-1}}{2}\,dx\,x^{\frac{d}{2}-1}$:
$\mathbb{B}^{\rm UV}_{ij;kl}(0;\delta_{\rm IR})=\frac{\Omega_{d-1}}{4(2\pi)^{d}}\int_{0}^{\infty}dx\,x^{\frac{d}{2}-1}\,\Big\langle\sum_{a,b}\mathbf{P}_{a}(\hat{\mathbf{n}})_{ik}\,\mathbf{P}_{b}(\hat{\mathbf{n}})_{jl}\,\mathcal{I}_{ab}(x;\delta_{\rm IR})\Big\rangle_{\hat{\mathbf{n}}}$, (183)
@@ B.1 Weights $J_{11}$, $J_{22}$ and $J_{12}$ in the limit $\mathbf{Z}_{t}=\mathbf{1},\mathbf{Y}=0$
which can be expressed through Lauricella $F_{D}^{(3)}$ as quoted in the main text.
@@ B.2 Proof of the inequality $J_{12}^{2}\leq J_{11}J_{22}$
Average both sides over $S^{2}$ and use Cauchy–Schwarz:
@@ Appendix C Linear stability of the generalized Wilson–Fisher fixed point
The coupled fixed points exist iff $\Delta\in\mathbb{R}$, i.e. iff $\rho\geq\sqrt{3}/2$ (which is the same condition as $4J_{12}^{2}\geq 3J_{11}J_{22}$ in the original variables). For the IR–stable branch $S_{-}$ one finds that
$g_{1}^{\star}=g_{2}^{\star}=g^{\star},\qquad\lambda_{1}^{\star}=\frac{g^{\star}}{J_{11}},\quad\lambda_{2}^{\star}=\frac{g^{\star}}{J_{22}},\quad\lambda_{3}^{\star}=\frac{g_{3}^{\star}}{\sqrt{J_{11}J_{22}}}.$ (230)
@@
$g_{3}^{\star}=\frac{c}{2(4\rho^{2}+1)}\Big(4\rho-\Delta\Big),\qquad g^{\star}=\frac{c}{2(4\rho^{2}+1)}\Big(4\rho^{2}+3+2\rho\,\Delta\Big),$ (231)
@@
$(\lambda_{1}^{\star},\lambda_{2}^{\star},\lambda_{3}^{\star})=\left(\frac{24\pi^{2}}{5}\varepsilon,\frac{24\pi^{2}}{5}\varepsilon,\frac{8\pi^{2}}{5}\varepsilon\right),$ (232)
@@
Let $t\equiv\ln\mu$. Linearizing the RG flow around a fixed point $g_{a}^{\star}$ gives
$\frac{d}{dt}\,\delta g_{a}=\sum_{b}\mathbb{M}_{ab}\,\delta g_{b},\qquad\mathbb{M}_{ab}\equiv\left.\frac{\partial\beta_{g_{a}}}{\partial g_{b}}\right|_{g=g^{\star}}.$ (233)
With our convention $\beta=\mu\frac{d}{d\mu}$, the IR limit corresponds to $t\to$-$\infty$. Thus the fixed point is IR–attractive iff all eigenvalues $\omega$ of $\mathbb{M}$ are positive, because $\delta g\sim e^{\omega t}$ then decays as $t\to$-$\infty$.
@@
$\mathbb{M}=\begin{pmatrix}-\varepsilon+6ag_{1}&0&6ag_{3}\\ 0&-\varepsilon+6ag_{2}&6ag_{3}\\ ag_{3}&ag_{3}&-\varepsilon+a(g_{1}+g_{2})+8a\rho g_{3}\end{pmatrix}_{g=g^{\star}}.$ (234)
At the coupled fixed points with $g_{1}^{\star}=g_{2}^{\star}=g^{\star}$ and $g_{3}^{\star}\neq 0$ one may simplify the diagonal entries to obtain
$\mathbb{M}=\begin{pmatrix}4a(g^{\star}-\rho g_{3}^{\star})&0&6ag_{3}^{\star}\\ 0&4a(g^{\star}-\rho g_{3}^{\star})&6ag_{3}^{\star}\\ ag_{3}^{\star}&ag_{3}^{\star}&4a\rho g_{3}^{\star}\end{pmatrix}$. (235)
@@
so that $\delta g_{-}$ is the “antisymmetric” deformation (splitting $g_{1}$ and $g_{2}$), while $\delta g_{+}$ is the “symmetric” deformation (changing them equally). From (235) one immediately sees that $\delta g_{-}$ decouples because the third row contains $\delta g_{1}+\delta g_{2}$ only. Explicitly,
$\frac{d}{dt}\delta g_{-}=\omega_{-}^{(\text{as})}\,\delta g_{-},\qquad\omega_{-}^{(\text{as})}=4a\big(g^{\star}-\rho g_{3}^{\star}\big).$ (237)
@@
$\frac{d}{dt}\begin{pmatrix}\delta g_{+}\\ \delta g_{3}\end{pmatrix}=\begin{pmatrix}4a(g^{\star}-\rho g_{3}^{\star})&6ag_{3}^{\star}\\ 2ag_{3}^{\star}&4a\rho g_{3}^{\star}\end{pmatrix}\begin{pmatrix}\delta g_{+}\\ \delta g_{3}\end{pmatrix}$. (239)
@@
Still at $h_{i}=0$, consider infinitesimal parity–breaking couplings $\lambda_{4,5}$ around the parity–symmetric fixed point $(\lambda_{4}^{\star},\lambda_{5}^{\star})=(0,0)$. Linearizing the beta functions from Sec. 5 gives
$\frac{d}{dt}\begin{pmatrix}\lambda_{4}\\ \lambda_{5}\end{pmatrix}=\mathbb{M}_{45}\begin{pmatrix}\lambda_{4}\\ \lambda_{5}\end{pmatrix},\qquad\mathbb{M}_{45}=-\varepsilon\,\mathbf{1}+\frac{3}{16\pi^{2}}\begin{pmatrix}J_{11}\lambda_{1}^{\star}+2J_{12}\lambda_{3}^{\star}&J_{22}\lambda_{3}^{\star}\\ J_{11}\lambda_{3}^{\star}&J_{22}\lambda_{2}^{\star}+2J_{12}\lambda_{3}^{\star}\end{pmatrix}$. (244)
Using $g_{1}^{\star}=g_{2}^{\star}=g^{\star}$ and $g_{3}^{\star}\neq 0$ one finds that the diagonal entries coincide and simplify to $\varepsilon/2$, while the off–diagonal product yields $\sqrt{(\mathbb{M}_{45})_{12}(\mathbb{M}_{45})_{21}}=3ag_{3}^{\star}$. Therefore the two eigenvalues in the $(\lambda_{4},\lambda_{5})$ sector are
$\omega_{4,5}^{(\pm)}=\frac{\varepsilon}{2}\pm 3a\,g_{3}^{\star}=\frac{\varepsilon}{2}\left(1\pm\frac{4\rho-\sqrt{4\rho^{2}-3}}{4\rho^{2}+1}\right).$ (245)
@@
In this appendix we analyze the linear stability of the fully interacting sector in the $Z_{t}=1,\;Y=0$ limit, where all quartic couplings $\lambda_{1,\dots,5}$ are allowed to be nonzero. The one–loop beta functions read
@@
In this appendix we analyze the linearized RG flow of the cubic couplings $h_{i}$ and the quadratic sector around the fully coupled fixed line found in the quartic sector (defined by $Z_{t}=\mathbf{1}$ and $Y=0$). Throughout we work in $D=4-\varepsilon$.
@@
$c\;\equiv\;\frac{16\pi^{2}}{3}\,\varepsilon,\qquad u_{i}\;\equiv\;\frac{g_{i}}{c}\,.$ (261)
@@
In the $Z_{t}=\mathbf{1}$, $Y=0$ limit, the one-loop beta functions for the cubic couplings $h_{1},h_{2},h_{3},h_{4}$ read
$\beta_{h_{1}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{1}+\frac{3}{16\pi^{2}}\Bigl(J_{11}\lambda_{1}h_{1}+2J_{12}\lambda_{4}h_{3}+J_{22}\lambda_{3}h_{4}\Bigr),$ (265)
$\beta_{h_{2}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{2}+\frac{3}{16\pi^{2}}\Bigl(J_{22}\lambda_{2}h_{2}+2J_{12}\lambda_{5}h_{4}+J_{11}\lambda_{3}h_{3}\Bigr),$ (266)
$\beta_{h_{3}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{3}$
@@
$\beta_{h_{4}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{4}$
@@
$\beta_{\bm{h}}\;=\;-\Bigl(1+\frac{\varepsilon}{2}\Bigr)\bm{h}+\frac{1}{16\pi^{2}}\,\mathbb{A}_{h}(\lambda,J)\,\bm{h},\qquad\bm{h}\equiv(h_{1},h_{2},h_{3},h_{4})^{\sf T}.$ (269)
@@
$\omega_{h}(x)=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)+\frac{\varepsilon}{3}\,x\qquad(x\in\{0,0,1,3\}).$ (280)
@@
All are negative for small $\varepsilon>0$, so any nonzero cubic couplings are relevant perturbations of the quartic fixed line in $4-\varepsilon$ dimensions. In particular, there is no perturbative fixed point with $h_{i}\neq 0$ controlled by the $4-\varepsilon$ expansion: the equation $\beta_{\bm{H}}=0$ admits only $\bm{H}=0$ at this order.
@@
The one-loop RG equations in the $Z_{t}=\mathbf{1}$, $Y=0$ limit can be written as
@@
$\beta_{\bm{\mathcal{m}}}=-2\,\bm{\mathcal{m}}-\frac{\varepsilon}{3}\,\mathbb{M}_{m}(u,\rho)\,\bm{\mathcal{m}}\qquad\text{(at linear order in $m^{2}$ and with $h=0$)},$ (285)
@@
$\omega_{m}(x)=-2-\frac{\varepsilon}{3}x,\qquad x\in\{0,0,1\},$ (290)

### v2

@@ 5.2 Turning on cross-gradients: $\mathbf{Z}_{t}=\mathbf{1}$ and $\mathbf{Y}\neq 0$
so the kernel $\mathcal{K}$ is literally the (direction–independent) two–particle density of states per logarithmic interval in the RG scale, while the weights (5) are obtained by dressing this phase space by the constant coherence factors in the projectors $\Pi_{\pm}$.
@@ 6 Conclusion
Technically, we have shown that in dimensional regularization with minimal subtraction near $D=4-\epsilon$ all one–loop ultraviolet divergences are governed by two universal master objects: the tadpole with a single mass insertion and the zero–momentum bubble. Using a spectral decomposition of the anisotropic quadratic operator, we derived compact, basis–covariant expressions for their pole parts in terms of the direction–dependent eigenvalues and projectors of the kinetic matrix. Both master residues are dressed by a single positive kernel, which admits a transparent interpretation as a two–particle energy denominator, or equivalently as a two–particle phase–space weight per logarithmic interval in the RG scale. This representation makes the physical origin of all anisotropy–dependent coefficients in the beta functions explicit.
@@ A.1 Tadpole integral
where $\mathbf{K}$ is given by (31). After introduction of such regulator into (41) the original integral
@@
where $\mathbf{p}\equiv\rho\hat{\mathbf{n}}$ with $\rho\equiv|\mathbf{p}|\geq 0$, and $x\equiv\rho^{2};\Omega_{d-1}$ is the solid angle on the unit sphere and $\Omega_{d-1}=2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$. We also remind, that $\langle\ldots\rangle_{\hat{\mathbf{n}}}$ is given by (44). Hence, we arrive at the regulated representation
@@
Our aim is to regularize the divergences in the latter expression. To this end let us use Schwinger parametrization
@@
$\times\int_{0}^{\infty}t^{-1/2}e^{-t\delta_{\rm IR}^{2}}dt\;t^{-d/2}\,\sum_{a}\left\langle s_{a}(\hat{\mathbf{n}})^{-d/2}\,\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}(\hat{\mathbf{n}})\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}$, (164)
@@
$\mathbf{T}_{0}(\delta_{\rm IR})$ $=\frac{1}{2^{d+1}\pi^{d/2}\sqrt{\pi}}\int_{0}^{\infty}t^{-\frac{(d+1)}{2}}e^{-t\delta_{\rm IR}^{2}}dt\;\sum_{a}\left\langle s_{a}^{-\frac{d}{2}}\,\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}(\hat{\mathbf{n}})\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}$, (165)
where we use $\Omega_{d-1}=2\pi^{d/2}/\Gamma\left(\frac{d}{2}\right)$. Using $d=3-\epsilon$, the $L=0$ contribution $\mathbf{T}_{0}$ is a scaleless power-divergent integral and vanishes in pure dimensional regularization. With the auxiliary IR mass $\delta_{\rm IR}>0$ kept finite one finds $\mathbf{T}_{0}(\delta_{\rm IR})\propto\delta_{\rm IR}^{2}$ (and may contain a pole $\propto\delta_{\rm IR}^{2}/\epsilon$), hence it vanishes as $\delta_{\rm IR}\to 0$ and does not contribute to the MS counterterms of the original theory [^8] I.e. in $d=3-\epsilon$ we use $\int_{0}^{\infty}dtt^{-\frac{d+1}{2}}e^{-t\delta_{\mathrm{IR}}^{2}}=\left(\delta_{\mathrm{IR}}^{2}\right)^{\frac{d-1}{2}}\Gamma\left(\frac{1-d}{2}\right)=\left(\delta_{\mathrm{IR}}^{2}\right)^{1-\epsilon/2}\Gamma\left(-1+\frac{\epsilon}{2}\right)\sim-\frac{2}{\epsilon}\delta_{\mathrm{IR}}^{2}+\mathcal{O}\left(\delta_{\mathrm{IR}}^{2}\ln\delta_{\mathrm{IR}}^{2}\right)$.. Therefore the MS pole relevant for renormalization starts at $\mathcal{O}(\mathbf{L})$, i.e. comes from $\mathbf{T}_{1}$.
@@ A.2 Bubble integrals
$\mathbf{B}_{ij;kl}(k)=\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\mathbf{G}_{ik}(p)\mathbf{G}_{jl}(p+k)$
@@
$\mathbf{B}_{ij;kl}(0;\delta_{\rm IR})\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\,\mathbf{G}_{\delta,ik}(p)\,\mathbf{G}_{\delta,jl}(p),$ (176)
@@
$\mathbf{B}_{ij;kl}(0;\delta_{\rm IR})=\frac{1}{2}\int\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\sum_{a,b}\tilde{\mathbf{P}}_{a}(\mathbf{p})_{ik}\,\tilde{\mathbf{P}}_{b}(\mathbf{p})_{jl}\,\mathcal{K}\!\big(\kappa_{a}(\mathbf{p})+\delta_{\rm IR}^{2},\;\kappa_{b}(\mathbf{p})+\delta_{\rm IR}^{2}\big),$ (180)
@@
Switching to $x\equiv|\mathbf{p}|^{2}$ and use $d^{d}\mathbf{p}=\frac{\Omega_{d-1}}{2}\,dx\,x^{\frac{d}{2}-1}$:
$\mathbf{B}^{\rm UV}_{ij;kl}(0;\delta_{\rm IR})=\frac{\Omega_{d-1}}{4(2\pi)^{d}}\int_{0}^{\infty}dx\,x^{\frac{d}{2}-1}\,\Big\langle\sum_{a,b}\mathbf{P}_{a}(\hat{\mathbf{n}})_{ik}\,\mathbf{P}_{b}(\hat{\mathbf{n}})_{jl}\,\mathcal{I}_{ab}(x;\delta_{\rm IR})\Big\rangle_{\hat{\mathbf{n}}}$, (182)
@@
$\mathbf{B}_{ij;kl}(0)]_{\rm pole}=\frac{1}{4\pi^{2}\epsilon}\Big\langle\sum_{a,b}\mathcal{K}(s_{a},s_{b})\,\mathbf{P}_{a}(\hat{\mathbf{n}})_{ik}\,\mathbf{P}_{b}(\hat{\mathbf{n}})_{jl}\Big\rangle_{\hat{\mathbf{n}}}$, (187)
@@ B.1 Weights $J_{11}$, $J_{22}$ and $J_{12}$ in the limit $\mathbf{Z}_{t}=\mathbf{1},\mathbf{Y}=0$
and this could be expressed through Lauricella hypergeometric functions $F_{D}^{(3)}$ as quoted in the main text.
@@ B.2 Proof of the inequality $J_{12}^{2}\leq J_{11}J_{22}$
Average both sides over $S^{2}$ and use Cauchy–Bunyakovsky–Schwarz:
@@ Appendix C Linear stability of the generalized Wilson–Fisher fixed point
The coupled fixed points exist if and only if $\Delta\in\mathbb{R}$, i.e. iff $\rho\geq\sqrt{3}/2$ (which is the same condition as $4J_{12}^{2}\geq 3J_{11}J_{22}$ in the original variables). For the IR–stable branch $S_{-}$ one finds that
$g_{1}^{*}=g_{2}^{*}=g^{*},\qquad\lambda_{1}^{*}=\frac{g^{*}}{J_{11}},\quad\lambda_{2}^{*}=\frac{g^{*}}{J_{22}},\quad\lambda_{3}^{*}=\frac{g_{3}^{*}}{\sqrt{J_{11}J_{22}}}.$ (229)
@@
$g_{3}^{*}=\frac{c}{2(4\rho^{2}+1)}\Big(4\rho-\Delta\Big),\qquad g^{*}=\frac{c}{2(4\rho^{2}+1)}\Big(4\rho^{2}+3+2\rho\,\Delta\Big).$ (230)
@@
$(\lambda_{1}^{*},\lambda_{2}^{*},\lambda_{3}^{*})=\left(\frac{24\pi^{2}}{5}\varepsilon,\frac{24\pi^{2}}{5}\varepsilon,\frac{8\pi^{2}}{5}\varepsilon\right),$ (231)
@@
Let $t\equiv\ln\mu$. Linearizing the RG flow around a fixed point $g_{a}^{*}$ gives
$\frac{d}{dt}\,\delta g_{a}=\sum_{b}\mathbb{M}_{ab}\,\delta g_{b},\qquad\mathbb{M}_{ab}\equiv\left.\frac{\partial\beta_{g_{a}}}{\partial g_{b}}\right|_{g=g^{*}}.$ (232)
With our convention $\beta_{g}=\mu\frac{dg}{d\mu}$, the IR limit corresponds to $t\to$-$\infty$. Thus the fixed point is IR–attractive iff all eigenvalues $\omega$ of $\mathbb{M}$ are positive, because $\delta g\sim e^{\omega t}$ then decays as $t\to$-$\infty$.
@@
$\mathbb{M}=\begin{pmatrix}-\varepsilon+6ag_{1}&0&6ag_{3}\\ 0&-\varepsilon+6ag_{2}&6ag_{3}\\ ag_{3}&ag_{3}&-\varepsilon+a(g_{1}+g_{2})+8a\rho g_{3}\end{pmatrix}_{g=g^{*}}.$ (233)
At the coupled fixed points with $g_{1}^{*}=g_{2}^{*}=g^{*}$ and $g_{3}^{*}\neq 0$ one may simplify the diagonal entries to obtain
$\mathbb{M}=\begin{pmatrix}4a(g^{*}-\rho g_{3}^{*})&0&6ag_{3}^{*}\\ 0&4a(g^{*}-\rho g_{3}^{*})&6ag_{3}^{*}\\ ag_{3}^{*}&ag_{3}^{*}&4a\rho g_{3}^{*}\end{pmatrix}$. (234)
@@
so that $\delta g_{-}$ is the “antisymmetric” (short notation“(as)”) deformation (splitting $g_{1}$ and $g_{2}$), while $\delta g_{+}$ is the “symmetric” (short notation “(sym)”) deformation (changing them equally). From (234) one immediately sees that $\delta g_{-}$ decouples because the third row contains $\delta g_{1}+\delta g_{2}$ only. Explicitly,
$\frac{d}{dt}\delta g_{-}=\omega_{-}^{(\text{as})}\,\delta g_{-},\qquad\omega_{-}^{(\text{as})}=4a\big(g^{*}-\rho g_{3}^{*}\big).$ (236)
@@
$\frac{d}{dt}\begin{pmatrix}\delta g_{+}\\ \delta g_{3}\end{pmatrix}=\begin{pmatrix}4a(g^{*}-\rho g_{3}^{*})&6ag_{3}^{*}\\ 2ag_{3}^{*}&4a\rho g_{3}^{*}\end{pmatrix}\begin{pmatrix}\delta g_{+}\\ \delta g_{3}\end{pmatrix}$. (238)
@@
Still at $h_{i}=0$, consider infinitesimal parity–breaking couplings $\lambda_{4,5}$ around the parity–symmetric fixed point $(\lambda_{4}^{*},\lambda_{5}^{*})=(0,0)$. Linearizing the beta functions from Sec. 5 gives
$\frac{d}{dt}\begin{pmatrix}\lambda_{4}\\ \lambda_{5}\end{pmatrix}=\mathbb{M}_{45}\begin{pmatrix}\lambda_{4}\\ \lambda_{5}\end{pmatrix},\qquad\mathbb{M}_{45}=-\varepsilon\,\mathbf{1}+\frac{3}{16\pi^{2}}\begin{pmatrix}J_{11}\lambda_{1}^{*}+2J_{12}\lambda_{3}^{*}&J_{22}\lambda_{3}^{*}\\ J_{11}\lambda_{3}^{*}&J_{22}\lambda_{2}^{*}+2J_{12}\lambda_{3}^{*}\end{pmatrix}$. (243)
Using $g_{1}^{*}=g_{2}^{*}=g^{*}$ and $g_{3}^{*}\neq 0$ one finds that the diagonal entries coincide and simplify to $\varepsilon/2$, while the off–diagonal product yields $\sqrt{(\mathbb{M}_{45})_{12}(\mathbb{M}_{45})_{21}}=3ag_{3}^{*}$. Therefore the two eigenvalues in the $(\lambda_{4},\lambda_{5})$ sector are
$\omega_{4,5}^{(\pm)}=\frac{\varepsilon}{2}\pm 3a\,g_{3}^{*}=\frac{\varepsilon}{2}\left(1\pm\frac{4\rho-\sqrt{4\rho^{2}-3}}{4\rho^{2}+1}\right).$ (244)
@@
In this appendix we analyze the linear stability of the fully interacting sector in the $\mathbf{Z}_{t}=1,\;\mathbf{Y}=0$ limit, where all quartic couplings $\lambda_{1,\dots,5}$ are allowed to be nonzero. The one–loop beta functions read
@@
In this appendix we analyze the linearized RG flow of the cubic couplings $h_{i}$ and the quadratic sector around the fully coupled fixed line found in the quartic sector (defined by $\mathbf{Z}_{t}=\mathbf{1}$ and $\mathbf{Y}=0$). Throughout we work in $D=4-\epsilon$.
@@
$c\;\equiv\;\frac{16\pi^{2}}{3}\,\epsilon,\qquad u_{i}\;\equiv\;\frac{g_{i}}{c}\,.$ (260)
@@
In the $\mathbf{Z}_{t}=\mathbf{1}$, $\mathbf{Y}=0$ limit, the one-loop beta functions for the cubic couplings $h_{1},h_{2},h_{3},h_{4}$ read
$\beta_{h_{1}}=-\Bigl(1+\frac{\epsilon}{2}\Bigr)h_{1}+\frac{3}{16\pi^{2}}\Bigl(J_{11}\lambda_{1}h_{1}+2J_{12}\lambda_{4}h_{3}+J_{22}\lambda_{3}h_{4}\Bigr),$ (264)
$\beta_{h_{2}}=-\Bigl(1+\frac{\epsilon}{2}\Bigr)h_{2}+\frac{3}{16\pi^{2}}\Bigl(J_{22}\lambda_{2}h_{2}+2J_{12}\lambda_{5}h_{4}+J_{11}\lambda_{3}h_{3}\Bigr),$ (265)
$\beta_{h_{3}}=-\Bigl(1+\frac{\epsilon}{2}\Bigr)h_{3}$
@@
$\beta_{h_{4}}=-\Bigl(1+\frac{\epsilon}{2}\Bigr)h_{4}$
@@
$\beta_{\bm{h}}\;=\;-\Bigl(1+\frac{\epsilon}{2}\Bigr)\bm{h}+\frac{1}{16\pi^{2}}\,\mathbb{A}_{h}(\lambda,J)\,\bm{h},\qquad\bm{h}\equiv(h_{1},h_{2},h_{3},h_{4})^{\sf T}.$ (268)
@@
$\beta_{\bm{H}}\;=\;-\Bigl(1+\frac{\epsilon}{2}\Bigr)\bm{H}+\frac{\epsilon}{3}\,\mathbb{M}_{h}(u,\rho)\,\bm{H},$ (271)
@@
$\omega_{h}(x)=-\Bigl(1+\frac{\epsilon}{2}\Bigr)+\frac{\epsilon}{3}\,x,\qquad(x\in\{0,0,1,3\}).$ (279)
@@
$\omega_{h}=\Bigl\{-1-\frac{\epsilon}{2},\;-1-\frac{\epsilon}{2},\;-1-\frac{\epsilon}{6},\;-1+\frac{\epsilon}{2}\Bigr\}.$ (280)
All are negative for small $\epsilon>0$, so any nonzero cubic couplings are relevant perturbations of the quartic fixed line in $4-\epsilon$ dimensions. In particular, there is no perturbative fixed point with $h_{i}\neq 0$ controlled by the $4-\epsilon$ expansion: the equation $\beta_{\bm{H}}=0$ admits only $\bm{H}=0$ at this order.
@@
The one-loop RG equations in the $\mathbf{Z}_{t}=\mathbf{1}$, $\mathbf{Y}=0$ limit can be written as
@@
$\beta_{\bm{\mathcal{m}}}=-2\,\bm{\mathcal{m}}-\frac{\epsilon}{3}\,\mathbb{M}_{m}(u,\rho)\,\bm{\mathcal{m}}\qquad\text{(at linear order in $m^{2}$ and with $h=0$)},$ (284)
@@
$\omega_{m}(x)=-2-\frac{\epsilon}{3}x,\qquad x\in\{0,0,1\},$ (289)